OpenStudy (anonymous):

The temperature of water in a certain lake on a day in October can be determined by using the model y=15.2-0.537x where x is the number of feet down from the surface of the lake and y is the Celsius temperature of the water at that depth. Based on this model, how deep in the lake is the water 9 degrees? (Round to the nearest foot.) a. 12 c. 67 b. 45 d. 30

OpenStudy (anonymous):

Let Y = 9 because the y is the Celsius temperature of the water. y=15.3-0.537x 9=15.3-0.537x x=11.7318